# Wei Jiang    Jiayu Yang    Yongqi Zhai    Feng Gao    Ronggang Wang

### v9

@@ 4.4.3 On Forward MACs
Compared with ${}^{+}$, our ${}^{++}$, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM (Liu et al., 2023), our ${}^{++}$ demonstrates significantly reduced MACs. Compared with ELIC (He et al., 2022), STF (Zou et al., 2022), and WACNN (Zou et al., 2022), our ${}^{++}$ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 (Cheng et al., 2020) as the basis for the transform module in our proposed ${}^{++}$. While the context module in Cheng’20 is simpler than that of ELIC, STF, and WACNN, the overall MACs of Cheng’20 are higher due to its transform modules having higher MACs. Since our ${}^{++}$ employs a modification of analysis transform and synthesis transform of Cheng’20, it results in higher MACs compared to ELIC and WACNN. In addition, since our ${}^{++}$ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
@@ 4.5.5 Analysis of Inter-Slice Global Context Module
Figure 15: Rate-distortion performance comparisons between ${}^{++}$ and ${}^{{++}}$ with vanilla quadratic complexity global spatial context modules.
@@
Table 5: BD-RATE(%) of training using different patch size. The anchor is ${}^{++}$ trained on $256\times 256$ patches.
@@
Table 6: Parameters and Forward Macs of context modules on Kodak.
@@ 4.5.7 Comparisons with Model with Vanilla Global Spatial Context Modules
The comparison between ${}^{++}$ and ${}^{++}$ with vanilla quadratic complexity global spatial context modules is illustrated in Fig. 15. It is evident that modeling global spatial contexts using linear complexity attention mechanisms does not result in performance degradation when compared to the vanilla attention mechanism.
@@ 4.5.8 Analysis of Training with Large Patches
In our training strategy, we use $512\times 512$ patches to train ${}^{++}$ during the rest $0.8$M steps. We compare the differences in rate-distortion performance between different patch sizes $\{256\times 256,320\times 320,512\times 512\}$ during the rest $0.8$M steps in Table 5. Using large patches further boost the model performance. Using $256\times 256$ patches cannot fully exploit the performance of the model. The size of latent representation is $16\times 16$ if $256\times 256$ patches are adopted. $16\times 16$ latent representation is insufficient for model to learn long-range or global dependency as the resolutions of input images of the codec could be 2K or 4K. Moreover, checkerboard partition (He et al., 2021) is employed, which makes the attention map sparse. Therefore, it is required to adopt large patches for better performance. Considering the overhead of training and model performance, adopting $512\times 512$ patches is the best choice.
@@ 4.5.9 Comparisons among Different Context Modules
As illustrated in Table 4, the proposed global context modules $g_{gc,intra}$ and $g_{gc,inter}$ contribute most to performance, channel-wise context module $g_{ch}$ has the second highest contribution to performance, and local sptail context module $g_{lc,attn}$ has the lowest contribution to performance. Since each context module has a different role to perform, it is imperative that they work together for performance enhancement. The complexity of each context module is presented in Table 6. The channel-wise context module has the most parameters and highest MACs. However, the MACs of the channel context module are only $1.77\%$ of the total MACs. The total MACs of all context modules are only $3.63\%$ of the total MACs.
@@ 5 Conclusion
In this paper, we propose a novel approach for capturing local spatial context using checkerboard attention, as well as linear complexity intra-slice and inter-slice global context modules, which significantly enhance the performance of the model while maintaining an acceptable linear complexity. Based on proposed context modules, we propose linear complexity multi-reference entropy model ${}^{++}$. Building upon ${}^{++}$, we obtain state-of-the-art model ${}^{++}$. ${}^{++}$ exhibits linear GPU memory consumption with resolution, making it highly suitable for high-resolution image coding. To make our ${}^{++}$ more practical, our future work will focus on investigating the asymmetrical design (Yang & Mandt, 2023) between the analysis and synthesis transforms, as well as lighter linear complexity multi-reference entropy model.

### v10

@@ 4.4.3 On Forward MACs
Compared with MLIC+, our MLIC++, which employs the proposed linear complexity global spatial context modules, has lower MACs. Compared with the recent LIC-TCM (Liu et al., 2023), our MLIC++ demonstrates significantly reduced MACs. Compared with ELIC (He et al., 2022), STF (Zou et al., 2022), and WACNN (Zou et al., 2022), our MLIC++ has higher MACs when the input is high-resolution image. One contributing factor to the higher MACs is the utilization of Cheng’20 (Cheng et al., 2020) as the basis for the transform module in our proposed MLIC++. While the context module in Cheng’20 is simpler than that of ELIC, STF, and WACNN, the overall MACs of Cheng’20 are higher due to its transform modules having higher MACs. Since our MLIC++ employs a modification of analysis transform and synthesis transform of Cheng’20, it results in higher MACs compared to ELIC and WACNN. In addition, since our MLIC++ employed an more advanced entropy model, which leads to higher MACs. However, the entropy model has modest effect on the overall MACs since the input image is down-sampled for four times in analysis transform, which implies that designing more advanced entropy models is more resource-efficient.
@@ 4.5.5 Analysis of Inter-Slice Global Context Module
Figure 15: Rate-distortion performance comparisons between MLIC++ and MLIC++ with vanilla quadratic complexity global spatial context modules.
@@
Table 5: BD-RATE(%) of training using different patch size. The anchor is MLIC++ trained on $256\times 256$ patches.
@@
Inference Time (s) $0.0036$ $0.0094$ $0.0179$ $0.0130$
Table 6: Parameters, forward Macs and inference time of context modules on Kodak.
@@ 4.5.7 Comparisons with Model with Vanilla Global Spatial Context Modules
The comparison between MLIC++ and MLIC++ with vanilla quadratic complexity global spatial context modules is illustrated in Fig. 15. It is evident that modeling global spatial contexts using linear complexity attention mechanisms does not result in performance degradation when compared to the vanilla attention mechanism.
@@ 4.5.8 Analysis of Training with Large Patches
In our training strategy, we use $512\times 512$ patches to train MLIC++ during the rest $0.8$M steps. We compare the differences in rate-distortion performance between different patch sizes $\{256\times 256,320\times 320,512\times 512\}$ during the rest $0.8$M steps in Table 5. Using large patches further boost the model performance. Using $256\times 256$ patches cannot fully exploit the performance of the model. The size of latent representation is $16\times 16$ if $256\times 256$ patches are adopted. $16\times 16$ latent representation is insufficient for model to learn long-range or global dependency as the resolutions of input images of the codec could be 2K or 4K. Moreover, checkerboard partition (He et al., 2021) is employed, which makes the attention map sparse. Therefore, it is required to adopt large patches for better performance. Considering the overhead of training and model performance, adopting $512\times 512$ patches is the best choice.
@@ 4.5.9 Comparisons among Different Context Modules
As illustrated in Table 4, the proposed global context modules $g_{gc,intra}$ and $g_{gc,inter}$ contribute most to performance, channel-wise context module $g_{ch}$ has the second highest contribution to performance, and local sptail context module $g_{lc,attn}$ has the lowest contribution to performance. Since each context module has a different role to perform, it is imperative that they work together for performance enhancement. The complexity of each context module is presented in Table 6. The channel-wise context module has the most parameters and highest MACs. However, the MACs of the channel context module are only $1.77\%$ of the total MACs. The total MACs of all context modules are only $3.63\%$ of the total MACs. Although $g_{ch}$ has high macs, its inference is fast due to the $3\times 3$ convolution being highly optimized on GPUs with CUDA/cuDNN [^3] https://github.com/pytorch/pytorch/blob/main/aten/src/ATen/native/Convolution.cpp while the window attention in $g_{lc,attn}$ and linear attention $g_{gc,intra}$ and $g_{gc,inter}$ are implemented in pure Python/Pytorch [^4] https://github.com/JiangWeibeta/MLIC/blob/main/MLIC%2B%2B/modules/transform/context.py. The $g_{gc,intra}$ consumes most time during inference could be attributed the partition of anchor part and non-anchor part.
@@ 5 Conclusion
In this paper, we propose a novel approach for capturing local spatial context using checkerboard attention, as well as linear complexity intra-slice and inter-slice global context modules, which significantly enhance the performance of the model while maintaining an acceptable linear complexity. Based on proposed context modules, we propose linear complexity multi-reference entropy model MEM++. Building upon MEM++, we obtain state-of-the-art model MLIC++. MLIC++ exhibits linear GPU memory consumption with resolution, making it highly suitable for high-resolution image coding. To make our MLIC++ more practical, our future work will focus on investigating the asymmetrical design (Yang & Mandt, 2023) between the analysis and synthesis transforms, as well as lighter linear complexity multi-reference entropy model.
Acknowledgement
This work is financially supported by Guangdong Provincial Key Laboratory of Ultra High Definition Immersive Media Technology (Grant No. 2024B1212010006), National Natural Science Foundation of China U21B2012, Shenzhen Science and Technology Program-Shenzhen Cultivation of Excellent Scientific and Technological Innovation Talents project (Grant No. RCJC20200714114435057), this work is also financially supported for Outstanding Talents Training Fund in Shenzhen.